# Geometry-Informed Neural Networks

### v3

@@ C.1 Metrics
and, similary, for the two-sided chamfer divergence
Reusing the notation from the paper, let $\mathcal{E}$ be the design region, $\delta\mathcal{E}$ the boundary of the design region, $\mathcal{I}$ the interface consisting of $n_{\mathcal{I}}$ connected components, $\mathcal{X}$ the domain, $\Omega$ the shape and $\delta\Omega$ its boundary.
Shape in design region.
We introduce two metrics to quantify how well a shape fits the design region. Intuitively for 3D, the first metric quantifies how much volume is outside the design region $\mathcal{E}$ compared to the overall volume that is available. The second metric compares how much surface area intersects the boundary of the design region.
@@
Fit to the interface.
To measure the goodness of fit to the interface, we use the one-sided chamfer distance of the boundary of the shape to the interface, as we do not care if some parts of the shape boundary are far away from the interface, as long as there are some parts of the shape which are close to the interface. A good fit is indicated by a 00 value.
@@
For the connectedness, we care whether the shape and whether the interfaces are connected. Since it is possible that the shape connects though paths that are outside the design region, we also introduce a metric that excludes such parts. The function $DC(\Omega)$ denotes all connected components of a shape $\Omega$ except the largest. We define the metrics as follows:
@@
  • $\frac{\operatorname{vol}(DC(\Omega))}{\operatorname{vol}(\mathcal{E})}$: To measure the $d$-volume (i.e. volume for $d=3$ and area for $d=2$) of disconnected components, we compute their volume and normalize it by the volume of the design region.
  • $\frac{\operatorname{vol}(DC(\Omega\cap\mathcal{E}))}{\operatorname{vol}(\mathcal{E})}$: Measures the $d$-volume of disconnected components inside the design region.
  • $\frac{CI(\Omega,\mathcal{I})}{n_{\mathcal{I}}}$ computes the share of connected interfaces. If an interface is an $\epsilon$-distance from a connected component of a shape, we consider it connected to the shape. This metric then represents the maximum number of connected interfaces of any connected component, divided by the total number of interface components. By default, we set $\epsilon=0.01$ when then domain bounds are comparable to the unit cube.
@@
There are many choices of smoothness measures in multiple dimensions. In this paper, we use a Monte Carlo estimate of the surface strain [61] (also mentioned in Section 4). To make the metric more robust to large outliers (e.g. tiny disconnected components have very large curvature and surface strain), we clip the surface strain of a sampled point $x_{i},i\in[N]$ with a value $\kappa_{\text{max}}=1000000$.
@@
C.2 Jet engine bracket
Eikonal loss facilitates learning.
Column 2 in Table 3 ablates the eikonal loss. Using the eikonal loss has a positive influence on the connectedness (see $b_{0}(\omega)$) and provides additional smoothness (see $E_{\text{strain}}(\Omega)$).
Connectedness loss is crucial for connected shapes.
Column 3 in Table 3 ablates the connectedness loss. Quantitatively, the zeroth Betti number $b_{0}(\Omega)$ (similary, $b_{0}(\Omega\cap\mathcal{E})$) is very high, i.e., there are many disconnected components. Furthermore, the share of connected interfaces $\frac{CI(\Omega,\mathcal{I})}{n_{\mathcal{I}}}$ is only $0.83$. Since for this problem there are 6 interfaces to connect, a value of $0.83$ implies that one of the interfaces is disconnected from the others.
Ablation eikonal connectedness smoothness log-smoothness base
smoothness $E_{\text{strain}}$ $E_{\text{strain}}$ $E_{\text{strain}}$ $E_{\text{log}}$ $E_{\text{strain}}$
$\lambda_{\text{eikonal}}$ 0 1 1 1 1
$\lambda_{\text{connectedness}}$ 1 0 1 1 1
$\lambda_{\text{smoothness}}$ 1 1 0 1 1
$\downarrow b_{0}(\Omega)$ 9 2 1 1 1
$\downarrow b_{0}(\Omega\cap E)$ 7 2 1 1 1
$\downarrow\frac{\operatorname{vol}(DC(\Omega))}{\operatorname{vol}(E)}$ 0.00 0.00 0.00 0.00 0.00
$\downarrow\frac{\operatorname{vol}(\Omega\setminus E)}{\operatorname{vol}(X\setminus E)}$ 0.00 0.00 0.00 0.00 0.00
$\uparrow\frac{CI(\Omega,I)}{n_{\mathcal{I}}}$ 0.67 0.83 1.00 1.00 1.00
$\downarrow CD_{1}(\Omega,I)$ 0.00 0.00 0.00 0.00 0.00
$\downarrow\frac{\operatorname{vol}(\Omega\cap\delta E)}{\operatorname{vol}(\delta E)}$ 0.00 0.00 0.00 0.00 0.00
$\downarrow\frac{\operatorname{vol}(DC(\Omega\cap E))}{\operatorname{vol}(E)}$ 0.00 0.00 0.00 0.00 0.00
$\downarrow E_{\text{strain}}(\Omega)$ 28,115 22,708 8,497 47,842 7,422
@@ C.2 Jet engine bracket
Table 3: Metrics for GINNS trained to produce a single shape of the jet engine bracket dataset.
Explicit smoothness also quantitatively improves smoothness.
Comparing Table 3, col. 4 to col. 6 shows that not using the smoothness loss leads to less smooth shapes. Qualitatively this is also depicted in Figure 3.
Explicit diversity loss improves diversity.
Comparing Table 4, col. 2 to col. 4 shows that not using the diversity loss halves the diversity $\delta_{\text{mean}}(S)$. Interestingly, also not using the eikonal loss reduces the diversity. We hypothesize, that the reason is that for training we compute a diversity loss on neural fields, sampled at points close to the individual boundaries. In contrast, the diversity metric (defined in section C.1) is computed using shapes at the zero level set of those fields with the chamfer-divergence as a pseudo-distance measure. Using the eikonal loss, leads to enforcing a more regular neural field, which in turn makes the diversity on neural fields more suitable.
Sampling generalizes better than fixed z.
Comparing Table 4, col. 3 to col. 4, the metrics for connectedness (e.g. $b_{0}(\Omega)$) and the number of connected interfaces improve when uniformly sampling the z from the domain during training.
Ablation diversity fixed z base
Training shapes 9 9 9
dim(z) 2 2 2
z sample method uniform fix uniform
smoothness $E_{\text{strain}}$ $E_{\text{strain}}$ $E_{\text{strain}}$
$\lambda_{\text{eikonal}}$ 1 1 1
$\lambda_{\text{connectedness}}$ 1 1 1
$\lambda_{\text{smoothness}}$ 1 1 1
$\lambda_{\text{div}}$ 0 1 1
$\downarrow b_{0}(\Omega)$ 1.00 3.00 1.11
$\downarrow b_{0}(\Omega\cap E)$ 1.00 3.00 1.11
$\downarrow\frac{\operatorname{vol}(DC(\Omega))}{\operatorname{vol}(E)}$ 0.00 0.00 0.00
$\downarrow\frac{\operatorname{vol}(\Omega\setminus E)}{\operatorname{vol}(X\setminus E)}$ 0.00 0.00 0.00
$\uparrow\frac{CI(\Omega,I)}{n_{\mathcal{I}}}$ 1.00 0.69 0.96
$\downarrow CD_{1}(\Omega,I)$ 0.00 0.01 0.00
$\downarrow\frac{\operatorname{vol}(\Omega\cap\delta E)}{\operatorname{vol}(\delta E)}$ 0.00 0.01 0.00
$\downarrow\frac{\operatorname{vol}(DC(\Omega\cap E))}{\operatorname{vol}(E)}$ 0.00 0.01 0.00
$\downarrow E_{\text{strain}}(\Omega)$ 23,059 27,563 31,592
$\uparrow\delta_{\text{mean}}$ 0.01 0.23 0.21
Table 4: Metrics for GINNS trained to produce multiple shapes of the jet engine bracket dataset. These are aggregated metrics averaged across all shapes.
@@ Appendix D Optimization
for a sequence $\{\mu_{0_{k}}\}$ with $\mu_{0_{k}}\leqslant\mu_{0_{k+1}}$ for all $k$ and $\mu_{0_{k}}\rightarrow\infty$. This so-called penalty method can however suffer from numerical instabilities for large $\mu_{0_{k}}$, hence the sequence is generally capped at a maximum value $\mu_{max}$. A further problem, which has recently been studied in regard to PINNs, is that the different objectives in 12 behave on different scales leading to instabilities in training as the gradients of the larger objective functions dominate training.
@@
Besides manual tuning of the weights $\mu_{i_{k}}$, several schemes to dynamically balance the different terms throughout training have been proposed, such as loss-balancing via the sub-gradients ([71]), via the eigenvalues of the Neural Tangent Kernel ([72]) or using a Soft Attention mechanism ([73]).
@@
$\min_{\theta}\max_{\lambda,\mu}\mathcal{L}(\theta,\lambda,\mu):=O(\theta)+\sum_{i=0}^{m}\lambda_{i}C_{i}(\theta)+\frac{1}{2}\mu_{0}\sum_{i=0}^{m}C_{i}^{2}(\theta)$ (14)
@@
$\max_{\lambda,\mu}\min_{\theta}\mathcal{L}(\theta,\lambda,\mu)\leqslant\min_{\theta}\max_{\lambda,\mu}\mathcal{L}(\theta,\lambda,\mu)$ (15)
@@
Note that this so-called dual update of the Lagrange multipliers is simply a gradient ascent step with learning rate $\mu_{0_{k}}$ for each multiplier $\lambda_{i_{k}}$. Typically, there is also an increase of $\mu_{0_{k}}$ up to maximum value $\mu_{max}$ as in the penalty method. Constrained optimization with neural networks using the ALM has been shown to perform well in previous works, such as in [74], [75], [76], [77] and [55].
In this classical ALM formulation, there is only a single penalty parameter $\mu_{0}$, which is monotonically increased during optimization. As outlined above, this is often insufficient to handle diverse constraints with different scales. Thus, we opt for the adaptive ALM proposed in [55] using adaptive penalty parameters for each constraint, solving 11 as the unconstrained optimization problem:
@@
In each epoch $k$ again a minimization step over the parameters $\theta_{k}$ via gradient descent is performed. Then the penalty parameters $\mu_{i_{k}}$, which are simultaneously the learning rate of the Lagrange multipliers $\lambda_{i_{k}}$, are updated using RMSprop followed by the gradient ascent step for $\lambda_{i_{k}}$
@@
Algorithm 1 shows the full algorithm used to train for $\mathcal{T}$ epochs and specifies the hyperparameters we used. The only difference to [55] is we set $\alpha=0.90$, which is the default value of RMSprop in PyTorch, instead of $\alpha=0.99$.
@@
5:     $\theta_{t}\leftarrow\operatornamewithlimits{argmin}_{\theta}\mathcal{L}(\theta_{t-1};\lambda_{t-1},\mu_{t-1})$triangleright primal update: a gradient descent step over $\theta$
@@
7:     $\mu_{t,i}\leftarrow\frac{\gamma}{\sqrt{\overline{v}_{t,i}}+\epsilon}$ $\forall i$triangleright penalty update
8:     $\lambda_{t,i}\leftarrow\lambda_{t-1,i}+\mu_{t,i}C_{i}(\theta_{t})$ $\forall i$triangleright dual update

### v4

@@ C.1 Metrics
We introduce several metrics for each constraint. We will use the chamfer divergence (Nguyen et al., 2021) to compute the divergence measure between two shapes $P$ and $Q$. For better interpretability, we take the square root of the common definition of chamfer divergence
and, similarly, for the two-sided chamfer divergence.
@@
Design region.
We introduce two metrics to quantify how well a shape fits the design region. Intuitively, for 3D, the first metric quantifies how much volume is outside the design region $\mathcal{E}$ compared to the overall available volume. The second metric compares how much surface area intersects the boundary of the design region. For both, the optimal values are 00, with lower being better.
@@
Interface.
To measure the goodness of fit to the interface, we use the one-sided chamfer distance of the boundary of the shape to the interface, as we do not care if some parts of the shape boundary are far away from the interface, as long as there are some parts of the shape which are close to the interface. The best fit is 00, with lower being better.
@@
For the connectedness, we care whether the shape itself and the interfaces are connected. Since the shape could potentially connect through paths outside the design region, we also introduce a connectedness metric that accounts for this undesirable effect. $DC(\Omega)$ denotes all connected components of a shape $\Omega$ except the largest. We define the metrics as follows:
@@
  • $\frac{\operatorname{vol}(DC(\Omega))}{\operatorname{vol}(\mathcal{E})}$: The normalized $d$-volume (i.e. volume for $d=3$ and area for $d=2$) of disconnected components.
  • $\frac{\operatorname{vol}(DC(\Omega\cap\mathcal{E}))}{\operatorname{vol}(\mathcal{E})}$: The normalized $d$-volume of disconnected components inside the design region.
  • $\frac{CI(\Omega,\mathcal{I})}{n_{\mathcal{I}}}$: The share of connected interfaces. If an interface is an $\epsilon$-distance from a connected component of a shape, we consider it connected to the shape. This metric then represents the maximum number of connected interfaces of any connected component, divided by the total number of interface components. By default, we set $\epsilon=0.01$ when the domain bounds are comparable to the unit cube.
@@
There are many choices of smoothness measures in multiple dimensions. In this paper, we use a Monte Carlo estimate of the surface strain (Goldman, 2005) (also mentioned in Section 4). To make the metric more robust to large outliers (e.g., tiny disconnected components have very large curvature and surface strain), we clip the surface strain of a sampled point $x_{i},i\in[N]$ with a value $\kappa_{\text{max}}=1,000,000$.
@@
C.2 Ablations
Using the metrics defined above, the impact of ablating each constraint, as well as ALM, is reported for the obstacle (Table 3), wheel (Table 4), and bracket (Table 5) tasks.
Table 3: Metrics for the ablations of the obstacle problem for multiple shape solutions.
Connectedness Interface Design region Diversity
Ablation $\downarrow b_{0}(\Omega)$ $\downarrow b_{0}(\Omega\cap E)$ $\downarrow\frac{\mathrm{vol}(DC(\Omega))}{\mathrm{vol}(E)}$ $\downarrow\frac{\mathrm{vol}(\Omega\setminus E)}{\mathrm{vol}(X\setminus E)}$ $\uparrow\frac{CI(\Omega,I)}{n_{\mathcal{I}}}$ $\downarrow CD_{1}(\Omega,I)$ $\downarrow\frac{\mathrm{vol}(\Omega\cap\delta E)}{\mathrm{vol}(\delta E)}$ $\uparrow\delta_{\text{mean}}$
(None) 1.11 1.11 0.02 0.00 0.39 0.03 0.01 0.15
Interface 0.00 1.00 0.00 0.00 0.00 0.00 0.00 0.00
Design region 1.00 1.00 0.00 0.53 2.00 0.00 0.71 0.08
Prescribed normal 0.00 1.00 0.00 0.00 0.00 0.00 0.00 0.00
Connectedness 0.56 1.00 0.00 0.00 0.00 0.00 0.00 1.45
Diversity 1.22 1.22 0.02 0.00 0.15 0.03 0.00 0.12
Surface strain 1.11 1.11 0.01 0.00 0.22 0.03 0.00 0.15
Augmented Lagrangian 1.22 1.22 0.02 0.00 0.28 0.03 0.01 0.14
Table 4: Metrics for the ablations of the wheel problem for multiple shape solutions.
Connectedness Interface Design region Diversity
Ablation $\downarrow b_{0}(\Omega)$ $\downarrow b_{0}(\Omega\cap E)$ $\downarrow\frac{\mathrm{vol}(DC(\Omega))}{\mathrm{vol}(E)}$ $\downarrow\frac{\mathrm{vol}(\Omega\setminus E)}{\mathrm{vol}(X\setminus E)}$ $\uparrow\frac{CI(\Omega,I)}{n_{\mathcal{I}}}$ $\downarrow CD_{1}(\Omega,I)$ $\downarrow\frac{\mathrm{vol}(\Omega\cap\delta E)}{\mathrm{vol}(\delta E)}$ $\uparrow\delta_{\text{mean}}$
(None) 1.11 1.00 0.00 0.00 1.00 0.00 0.07 0.31
Interface 1.00 1.00 0.00 0.00 0.44 0.07 0.02 1.10
Design region 6.89 1.00 0.06 0.11 1.00 0.00 0.44 0.68
Prescribed normal 1.00 1.00 0.00 0.00 1.00 0.00 0.03 0.19
Connectedness 1.22 1.22 0.02 0.00 0.89 0.00 0.11 0.34
Diversity 1.00 1.00 0.00 0.00 1.00 0.00 0.08 0.03
Rotational symmetry 1.00 1.00 0.00 0.00 1.00 0.00 0.06 0.54
Minimum thickness 1.11 1.00 0.00 0.00 1.00 0.00 0.10 0.38
Eikonal 1.00 1.00 0.00 0.00 1.00 0.00 0.09 0.03
Curvature & Diversity 1.00 1.00 0.00 0.00 1.00 0.00 0.08 0.03
Augmented Lagrangian 1.00 1.00 0.00 0.00 1.00 0.00 0.01 0.21
@@ C.2 Ablations
Table 5: Metrics for ablations of the jet engine bracket problem for single and multiple shape solutions.
Connectedness Interface Design region Smoothness Diversity
Ablation $\downarrow b_{0}(\Omega)$ $\downarrow b_{0}(\Omega\cap E)$ $\downarrow\frac{\mathrm{vol}(DC(\Omega))}{\mathrm{vol}(E)}$ $\downarrow\frac{\mathrm{vol}(\Omega\setminus E)}{\mathrm{vol}(X\setminus E)}$ $\uparrow\frac{CI(\Omega,I)}{n_{\mathcal{I}}}$ $\downarrow CD_{1}(\Omega,I)$ $\downarrow\frac{\mathrm{vol}(\Omega\cap\delta E)}{\mathrm{vol}(\delta E)}$ $\downarrow E_{\text{strain}}(\Omega)$ $\uparrow\delta_{\text{mean}}$
Multiple shapes
(None) 1.00 1.00 0.00 0.00 1.00 0.00 0.00 410 0.14
Interface 1.11 1.11 0.00 0.00 1.00 0.02 0.00 373 0.13
Design region 4.44 1.00 -0.01 0.97 1.00 0.00 0.14 95 0.09
Prescribed normal 1.00 1.00 0.00 0.00 1.00 0.01 0.00 343 0.13
Connectedness 25.11 19.44 0.00 0.00 0.17 0.01 0.00 178740 0.02
Diversity 1.00 1.00 0.00 0.00 1.00 0.00 0.00 341 0.05
Surface strain 1.00 1.00 0.00 0.00 1.00 0.00 0.00 263 0.10
Minimum thickness 1.00 1.00 0.00 0.00 1.00 0.01 0.00 432 0.15
Eikonal 6.67 6.44 0.00 0.00 1.00 0.00 0.00 722 0.07
Curvature & Diversity 1.00 1.00 0.00 0.00 1.00 0.00 0.00 305 0.05
Augmented Lagrangian 3.22 2.00 0.00 0.00 1.00 0.00 0.01 198 0.07
Single shape
(None) 1.00 1.00 0.00 0.00 1.00 0.01 0.00 291
Augmented Lagrangian 1.00 1.00 0.00 0.01 1.00 0.01 0.06 170
C.3 Runtimes
In addition to quantifying the impact of the constraint ablations on the metrics, we also report the iteration times in Table 6, focusing on the 3D jet engine bracket problem, which has the highest number and complexity of constraints. The other experiments show similar behavior. Most constraints have little effect on the time per iteration. Of the total runtime, the surface strain takes 10% (due to the Hessian) and the PH solver 75% (expensive multi-processed CPU task). The runtime increases when ablating the eikonal constraint as it destroys the geometric regularity of the implicit function, hindering efficient surface point sampling that usually takes 15% of total time. We use an A100-SXM GPU, and the peak memory usage for a batch of 9 shapes does not exceed 16 GB. These two losses also have a strong impact on the iterations needed. As discussed in the main text, adding the smoothness loss increases the number of iterations roughly two-fold, and adding the diversity increases it roughly five-fold. Overall, the biggest effect is not so much the number of constraints, but rather losses that are ill-conditioned. This conditioning issue is also known in the PINN literature, and its mitigation is an open and active research topic.
Table 6: Time per iteration when ablating each constraint. Largest differences from the baseline (None) indicate computationally expensive components of the method, most notably the connectedness constraint (due to the CPU-based PH calculation) and the boundary sampler required for calculating both the diversity and surface-strain.
Ablated constraint Time [ms] per iteration
(None) 260
Eikonal 291
Interface 264
Design region 262
Prescribed normal 265
Diversity 265
Surface strain 230
Surface strain & Diversity (no boundary point sampler) 187
Connectedness 94
C.4 Baselines
While a complete and fair comparison to a baseline is not available, we compare to two at least partially fitting baselines.
Topology optimization.
We consider classical TO, specifically, FeniTop (Jia et al., 2024) which implements the standard SIMP method with a popular FEM solver. We define a TO problem that is as similar as possible, applying a diagonal force to the top cylindrical pin interface and allowing a 7% volume fraction in the same design region. The other interfaces are fixed in the same way. The shape compliance is minimized for 400 iterations on a $104\times 172\times 60$ FEM grid (taking 190 min on a 32 core CPU to give a sense of runtime, although a fair timing comparison requires a more nuanced discussion). The produced shape is visualized in Figure 10.
We then compute the surface strain (the objective we use) for this TO shape and, conversely, the compliance for a GINN shape (Section 4.2; illustrated in the penultimate column of Figure 3). Unsurprisingly, both shapes perform best at the objective they are optimized for while satisfying the constraints up to the relevant precision. This serves as a sanity check and confirmation of the constraint satisfaction.
Figure 10: A solution to the jet engine bracket problem found by the FeniTop (Jia et al., 2024) topology optimization software. The quantitative metrics are reported in Table 7.
Table 7: GINN compared to the topology optimization baseline for a single jet engine bracket shape. Both methods perform best on the objectives they are optimized for while satisfying the shared constraints up to the relevant precision.
Metric Topology optimization GINN
$\downarrow$ Connectedness (0-th Betti-number) 1 1
$\downarrow$ Interface (Chamfer distance) 0.00 0.00
$\downarrow$ Design region (Volume outside) 0.00 0.00
$\downarrow$ Curvature 442 144
$\downarrow$ Compliance 0.99 0.344
Human-expert dataset.
A unique aspect of GINN is the data-free shape generative aspect. Comparison to classical TO is trivial since it is inherently limited to a single solution with null diversity. Instead, we use the simJEB (Whalen et al., 2021) dataset to give an intuitive estimate of the diversity of the produced results. The dataset is due to the design challenge on a related problem described in Section 4.2. The shapes in the dataset were produced by human experts, many of whom also used topology optimization. To compute the diversity metric, we sample 196 clean shapes from the simJEB dataset, producing a diversity of 0.099, and $14\times 14$ equidistant samples from the 2D latent space generative GINN model, producing a diversity of 0.167. Even though these sets are not directly comparable as they optimize for different objectives, these results indicate that GINNs can produce diversity on the same and larger magnitude as a dataset that required an estimated collective effort of 14 expert human years (Whalen et al., 2021).
@@ Appendix D Optimization
for a sequence $\{\mu_{0_{k}}\}$ with $\mu_{0_{k}}\leq\mu_{0_{k+1}}$ for all $k$ and $\mu_{0_{k}}\rightarrow\infty$. However, this penalty method can suffer from numerical instabilities for large $\mu_{0_{k}}$, hence the sequence is generally capped at a maximum value $\mu_{max}$. A further problem, which has recently been studied regarding PINNs, is that the different objectives in 12 behave on different scales, leading to instabilities in training as the gradients of the larger objective functions dominate training.
@@
Besides manual tuning of the weights $\mu_{i_{k}}$, several schemes to dynamically balance the different terms throughout training have been proposed, such as loss-balancing via the sub-gradients ((Wang et al., 2021)), via the eigenvalues of the neural tangent kernel (Wang et al., 2022) or using a soft-attention mechanism (McClenny & Braga-Neto, 2020).
@@
$\min_{\theta}\max_{\lambda,\mu}\mathcal{L}(\theta,\lambda,\mu):=O(\theta)+\sum_{i=0}^{m}\lambda_{i}C_{i}(\theta)+\frac{1}{2}\mu_{0}\sum_{i=0}^{m}C_{i}^{2}(\theta)\ .$ (14)
@@
$\max_{\lambda,\mu}\min_{\theta}\mathcal{L}(\theta,\lambda,\mu)\leq\min_{\theta}\max_{\lambda,\mu}\mathcal{L}(\theta,\lambda,\mu)$ (15)
@@
Note that this so-called dual update of the Lagrange multipliers is simply a gradient ascent step with learning rate $\mu_{0_{k}}$ for each multiplier $\lambda_{i_{k}}$. Typically, there is also an increase of $\mu_{0_{k}}$ up to maximum value $\mu_{max}$ as in the penalty method. Constrained optimization with neural networks using the ALM has been shown to perform well in previous works, such as in (Son et al., 2023), (Kotary & Fioretto, 2024), (Sangalli et al., 2021), (Fioretto et al., 2021), and (Basir & Senocak, 2023).
In this classical ALM formulation, there is only a single penalty parameter $\mu_{0}$, which is monotonically increased during optimization. As outlined above, this is often insufficient to handle diverse constraints with different scales. Thus, we opt for the adaptive ALM proposed in (Basir & Senocak, 2023) using adaptive penalty parameters for each constraint, solving 11 as the unconstrained optimization problem:
@@
In each epoch $k$, again a minimization step over the parameters $\theta_{k}$ via gradient descent is performed. Then the penalty parameters $\mu_{i_{k}}$, which are simultaneously the learning rate of the Lagrange multipliers $\lambda_{i_{k}}$, are updated using RMSprop followed by the gradient ascent step for $\lambda_{i_{k}}$
@@
Algorithm 1 shows the full algorithm used to train for $\mathcal{T}$ epochs and specifies the hyperparameters we used. The only difference to (Basir & Senocak, 2023) is that we set $\alpha=0.90$, which is the default value of RMSprop in PyTorch, instead of $\alpha=0.99$.
@@
5:     $\theta_{t}\leftarrow\operatornamewithlimits{argmin}_{\theta}\mathcal{L}(\theta_{t-1};\lambda_{t-1},\mu_{t-1})$primal update: a gradient descent step over $\theta$}
@@
7:     $\mu_{t,i}\leftarrow\frac{\gamma}{\sqrt{\overline{v}_{t,i}}+\epsilon}$ $\forall i$penalty update}
8:     $\lambda_{t,i}\leftarrow\lambda_{t-1,i}+\mu_{t,i}C_{i}(\theta_{t})$ $\forall i$dual update}